OpenStudy (anonymous):

can someone please help me with finding the equation of a best fit line in standard form? i think i have the concept but im sure im doing something wrong

OpenStudy (xanthe):

QN?

OpenStudy (anonymous):

heres the questions.....For problems 1 and 2, graph the data points. Draw the line of best fit and, using complete sentences, explain how you found it in 3-4 sentences. Write the equation for the line of best fit in standard form. (10, 150) (20, 275) (30, 340) (40, 400) (50, 475) Title: Intersection Cars Horizontal axis: Minutes elapsed Vertical axis: Total cars passing through IM currently working out number one and i have worked it out but believe i have error ed

OpenStudy (xanthe):

Have u drawn the graph?

OpenStudy (xanthe):

Use the equation y=mx+c

OpenStudy (anonymous):

yes i have and i see that the points i would have to use are (10, 150) and (50, 475)

OpenStudy (anonymous):

the equation im using is m = y2-y1 over x2-x1 to find the slope

OpenStudy (anonymous):

then the next equation im using is y - y1 = m (x - x1)

OpenStudy (xanthe):

Find m and c where m is the gradient and c is the y-intercept.The gradient of the points should be as far apart as possible and it shd be recommended to use the points from the best-fit line grapg that youve drawn and not the points given ie.:(10, 150) and (50, 475)

OpenStudy (anonymous):

for m I get 8.1 which im sure is wrong

OpenStudy (xanthe):